From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: July 23, 2009

Participants: Daniel, Mark I., Bill, Simon, Brian, Fernando, Mark S., Rich, Eugene

Next Meeting: July 30, 2009 @ 1:30 p.m.

General Construction Updates

 - Mark S. will be working exclusively on the solenoid for the next 5 days.
   We will have to absorb the manpower hit and deal with it.
 - The parts for the assembly jig on which the detector stack will be
   constructed will be put together as there is time.
 - Bill will take on the preparation for a stack assembly document. A
   draft will be prepared before assembly of the full-scale prototype and
   then fleshed out during the process with relevant photographs taken.
 - The composite frames for the two gas windows are in progress. So far the
   Rohacell core has been laminated to a G10 ring. One more G10 ring to go to
   complete the frame construction. The exposed Rohacell will then be sealed
   with epoxy. 
 - Mark S. prepared the PO for the cables that will be used to attach to 
   the cathodes for the external ground. He will also order the sealing 
   epoxy that is used to cover the conducting epoxy. The cables and associated
   pins should be here in a couple of weeks. The epoxy order has not yet
   been placed.
 - Rich will only be working with us for another couple of weeks before
   going back to UMass.
 - We need to be out of the clean room by the end of September. We are also
   thinking of carving out more space in EEL 126 for addition of a clean tent
   for construction.

Wire Frame Update

 - Brian repaired a third wire plane (called P4). Efforts were made to clean 
   the solder flux off the boards, but residue remains. Rich has done a 
   series of high voltage tests. One of the field wire busses on the HTVBs is 
   drawing sizeable current. After visual inspection for debris and cleaning 
   the PCB for a second and third time, the current on the supply dropped 
   from 50 uA to 15 uA. We are still devising better ways to clean the boards.
   Of  course the boards should have been cleaned at IUCF immediately after 
   the soldering was done. This will be mark of the Phase 3 discussion.
 - During inspection of the P4 wire frame, it was discovered that some of the
   components had the wrong values, some components were not properly
   soldered, some components were missing, and some were bad. A full check
   out has been done and Rich has prepared a list of components for repair.
   All other wire frames will also have to be rigorously checked.
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We will include
   the information on the wire electroplating process shortly.
 - The parts of the jig to use for the wire electroplating operation
   are in our hands. They will be assembled as we find time.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   when he gets a chance.

Cathode Update

 - Progress updates:
     * The edge sealing and gas hole inclusion has been completed on all
       of our cathode frames.
     * The 7th and last cathode has been tensioned and will be laminated
       to its frame this afternoon.
     * It was found that two of our 2 um Cu cathodes were made incorrectly
       with the center panel rotated by 180 deg. Bill has come up with a
       way to modify the frames and to modify a couple of cooling clamps to
       get around this. These two frames will have to be on the outside
       faces of the chamber. The layer with the two 5 um Cu boards will be
       in the middle of the prototype.
     * We will modify the layout of our cathode panel alignment fiducials
       to allow us to catch panel rotations. Details will be passed on to
     * One of our completed cathode boards is really contaminated with
       epoxy near the two seams. The epoxy manufacturer is sending us a
       solvent that they believe should allow the epoxy to be removed. It will
       arrive next week and we will carry out some tests.
     * The completed cathodes are sitting in a nitrogen environment to
       reduce the oxidation problems.
     * We have a bit of a back up doing the flatness scans of the completed
       cathodes. Mark S. has taught Brian how to complete the scans and a
       scan will be done this afternoon.
 - Bill stepped us through a work plan for the construction of the two
   cathode sandwiches for the prototype. This work will begin next week
   with the tensioning of the ground planes.
 - Bill is devising a work plan for the prototype construction that he
   will give to Chuck to put into a schedule for tracking.
 - The Fast Electronics group is handling the soldering of the rigid-flex
   assemblies. They now have all of the necessary components. The boards
   arrived poorly packaged and were noticeably warped. This has made the
   assembly tougher. Fernando came up with a scheme to deal with this, and
   the work to mount the components is underway. A QA plan for the finished
   boards needs to be put in place.
 - Bill has been working on the cathode construction document building in
   the pictures that are being taken. The current draft of the document 
   is located at /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. 
 - We are standing by waiting for information on the PR for the 2-um board
   material. We have learned that the order has gone out, but the Japanese
   company has still not indicated whether they will fill the order. A
   second company that can produce the board material has been identified.
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. Progress is slow and steady. The gas tubing is in 
   hand and will be laid out shortly.
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. We decided to have a separate meeting to discuss what work
   we would like to do on the small-scale prototype.
 - Beni has started a bench test plan for the full-scale prototype and
   is presently working on it after incorporating feedback. It was also
   suggested to send it to Curtis for feedback. He will post it as a 
   GlueX-note shortly.
 - The machined parts for the support/rotation of the full-scale prototype
   in the cosmic ray stand have been delivered. We are awaiting some purchased
   parts to arrive. After we have everything in hand, it will be assembled
   and test fitted.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.